OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all potential rational zeros f(x) = 14x^3 + 56x^2 + 2x - 7

OpenStudy (anonymous):

make all fractions \(\frac{p}{q}\) where \(p\) divides \(7\) and \(q\) divides \(14\) for example \[\pm1,\pm7,\pm\frac{1}{14}\]

OpenStudy (anonymous):

i left out one (well actually two if you include \(+\) and \(-\)

OpenStudy (anonymous):

Thank you! Which one did you leave out?

OpenStudy (anonymous):

lol the one that looks like \(\pm\frac{7}{14}\) better known as \(\pm\frac{1}{2}\)

OpenStudy (anonymous):

wouldnt it be 1/2 instead of 1/14?

OpenStudy (anonymous):

Oh thanks!